OpenStudy (anonymous):

Find the slope of the line through the pair of points. C(-6, -7), K(–1, -7) my answer; undefined ????

OpenStudy (anonymous):

(y2-y1)/(x2-x1)=(-7+7)/(-1+6) = 0

OpenStudy (a_clan):

0/5 =0

OpenStudy (anonymous):

If you drew a line through both points you'd see it's flat. It has no slope. So zero.